# Parameter-Efficient Continual Fine-Tuning: A Survey

### v1

@@ Abstract
The emergence of large pre-trained networks has revolutionized the AI field, unlocking new possibilities and achieving unprecedented performance. However, these models inherit a fundamental limitation from traditional Machine Learning approaches: their strong dependence on the i.i.d. assumption hinders their adaptability to dynamic learning scenarios. We believe the next breakthrough in AI lies in enabling efficient adaptation to evolving environments — such as the real world — where new data and tasks arrive sequentially. This challenge defines the field of Continual Learning (CL), a Machine Learning paradigm focused on developing lifelong learning neural models. One alternative to efficiently adapt these large-scale models is known Parameter-Efficient Fine-Tuning (PEFT). These methods tackle the issue of adapting the model to a particular data or scenario by performing small and efficient modifications, achieving similar performance to full fine-tuning. However, these techniques still lack the ability to adjust the model to multiple tasks continually, as they suffer from the issue of Catastrophic Forgetting. In this survey, we first provide an overview of CL algorithms and PEFT methods before reviewing the state-of-the-art on Parameter-Efficient Continual Fine-Tuning (PECFT). We examine various approaches, discuss evaluation metrics, and explore potential future research directions. Our goal is to highlight the synergy between CL and Parameter-Efficient Fine-Tuning, guide researchers in this field, and pave the way for novel future research directions.
@@ 1 Introduction
Deep learning models achieve impressive results by scaling their parameters and the data used during training. However, these models often require substantial resources, which makes it impractical to update them continuously. This limitation is especially significant when resources are constrained. Continual Learning (CL) provides a solution by enabling these models to learn and adapt to new information while retaining previously acquired knowledge. Specifically, CL methods address the challenge of training a single model to learn and retain knowledge across a sequence of tasks over time, ensuring that previously learned information is not forgotten.
CL is crucial for Large Pre-Trained Models (PTMs), which are expensive to train from scratch and can suffer from catastrophic forgetting during traditional fine-tuning for new tasks. Parameter-Efficient Fine-Tuning (PEFT) mitigate this by strategically adjusting only a small subset of parameters in the pre-trained model, significantly reducing computational costs while achieving similar performance. When combined with CL, PEFT allows PTMs to continuously learn and integrate new information while retaining past knowledge, in a resource-efficient manner. This paves the way for a new generation of large PTMs that can continuously evolve and adapt without extensive retraining.
The intersection between CL and PEFT is a fascinating area of study. On one hand, CL focuses on learning from dynamic and evolving data streams, allowing models to adapt to new tasks without forgetting previously learned knowledge. This is a critical aspect of machine learning, as it enables models to stay relevant in the face of changing data landscapes. On the other hand, PECFT enhances pre-trained models by strategically adjusting a limited subset of parameters during fine-tuning.
The combination of these two principles gives rise to Parameter Efficient Continual Fine Tuning (PECFT). PECFT combines the adaptability of CL with the efficiency of PEFT, enabling models to progressively adapt to new tasks while retaining past knowledge. This convergence introduces a new paradigm for large PTMs, allowing continuous evolution without extensive retraining. Works like Learning to Prompt [144] and Dual Prompt [143] have already explored this idea of CL and PEFT. It represents an exciting convergence of efficiency and adaptability, two traits that are increasingly recognized as critical in the fast-paced, ever-evolving world of machine learning. By balancing these elements, PECFT paves the way for the development of more effective, adaptable, and sustainable machine learning models that are capable of meeting the complex and growing demands of the field.
@@ 2.1.1 CL Challenges
Traditionally, machine learning models are usually trained on a finite, limited and static data set, limiting their adaptability. Naively training a model on new data to perform well on novel tasks can degrade its performance on previously learned tasks. This problem is known as Catastrophic Forgetting (CF) [101] and is the core challenge of CL. Theoretically, consequential re-training and adapting to new tasks can shift how features are represented within the model. This phenomenon, where the representation of features within a model evolves in a way that can negatively impact previous tasks, is known as Representation Drift [9, 54].
@@ 2.1.2 CL Scenarios
A second scenario is Class Incremental Learning (CIL), which, while similar to TIL in training, is more challenging in inference. This is because, during testing, the specific task information is unavailable, causing the model to have to predict among all observed classes, not only those of the task $t$. This lack of task information during testing requires the model to differentiate and classify all previously learned classes without additional context, making CIL significantly more challenging.
While TIL, CIL, DIL and OCL scenarios dominate much of the continual learning research, less frequently explored scenarios offer valuable insights into more realistic learning environments. Two such scenarios are Class-Incremental Learning with Repetition and Rainbow Memory.
@@ 3.1 Regularization-based CL
One alternative to naively training a model in a sequence of tasks is adding a constraint on the loss functions to encourage the mitigation of forgetting. This regularization can be imposed on weights by estimating the parameters importance so relevant weights do not drift significantly. One of the first approaches that proposed this was Elastic Weight Consolidation (EWC) [67], which captures the prior importance using a diagonally approximated Fisher information matrix. EWC is later improved by finding a better approximation of the fisher information matrix (sFIM).
Regularization can also be imposed on activations to prevent activation drift, which generally outperforms its weight-regularisation counterpart. One work in this line is Learning without Forgetting (LwF) [78], which prevents activations of the old network from drifting while learning new tasks. The less-forgetting learning penalizes the activation difference except for the fully-connected layer. Riemannian Walk [11] extends EWC by incorporating KL-divergence-based regularization, path integral methods to assess parameter importance and sample-based strategies for retaining past knowledge. Rotated EWC (R-EWC) [84] enhances EWC by reparameterizing the network to better align with the FIM, improving the diagonal approximation and reducing forgetting.
@@ 4 Parameter-Efficient Fine-Tuning
Table 2: Parameter-Efficient Fine-tuning Methods
Method Key Mechanism and Features
Adapter [A] Bottleneck layers ($2kd$/layer) with down/up projections. Effective task adaptation with minimal overhead
AdaptFormer [A] Specialized vision modules ($<2\%$) enhancing ViT transferability across tasks
Adapter Fusion [A] Two-stage fusion combining task adapters with knowledge distillation
Mera [A] Efficient integration of existing adapters using same-track combination strategy
AdaMix [A] Multiple parallel adapters providing diverse task perspectives
LoRA [R] Low-rank decomposition ($2dr$/layer) for weight updates. Zero inference overhead
QLoRA [R] 4-bit NormalFloat quantization + LoRA (0.1%) for memory-efficient training
DoRA [R] Separates magnitude and direction components for improved training dynamics
RoSA [R] Combines low-rank and sparse components for better outlier handling
VeRA [R] Shared frozen matrices with learnable scaling vectors
Prompt Tuning [A] Continuous trainable prefix tokens ($l\times d$) for task conditioning
P-Tuning [A] Hybrid discrete and continuous prompts with LSTM/MLP encoder
Prefix Tuning [A] Optimizable virtual tokens (0.1%) as prefix for generation
MPT [A] Distillation-based transferable prompts (0.035%) for multi-task
BitFit [S] Selective bias-term updates ($<0.1\%$) preserving model knowledge
Masking [S] Task-specific binary masks identifying crucial weights (3-10%)
[A]: Additive, [R]: Reparameterization, [S]: Selective
@@
PEFT strategies can be broadly categorized based on how they modify or add parameters to the pre-trained model. These categories include: (i) additive methods, which introduce new task-specific parameters to the model; (ii) reparameterization methods, which reparameterize the existing model parameters more efficiently; (iii) and selective methods, which fine-tune a subset of the model parameters based on their importance to the task. Since this paper focuses on the intersection of Continual Learning and Parameter-Efficient Fine-tuning, we will present an overview of these methods, introducing and discussing only a few of these — for a comprehensive comparison of different PEFT categories and methods, we refer readers to recent surveys by [160] and [147]. An summary of the PEFT methods we discuss can be seen in Table 2.
@@ 4.1.2 Prompt Based PEFT Methods
Building on this foundation, P-Tuning [85] was introduced to address a critical limitation — the inherent instability issues of discrete prompts. It introduces trainable continuous prompt embeddings that are concatenated with traditional discrete prompts. The core template structure is defined as:
@@
where $[P_{i}]$ represents continuous prompt embeddings that are optimized through backpropagation, while $x$ and $y$ represent the input and label respectively. The combination of trainable continuous embeddings with discrete prompts is done via a prompt encoder, i.e. LSTM or MLP, to model the dependencies between the embeddings. This hybrid approach significantly reduces performance variance, maintaining a consistent behavior where traditional methods showed up to 20% drops from single word changes in prompts.
@@
Taking prompt tuning to a multi-task setting, Multitask Prompt Tuning (MPT) [142] innovates by learning a single transferable prompt from multiple source tasks that can be adapted to target tasks. By operating in two stages — source training and target adaptation — MPT learns a task-shared prompt matrix through knowledge distillation. This prompt is decomposed into a task-shared matrix $P^{*}$ and low-rank task-specific matrices $W_{k}=u_{k}\otimes v_{k}^{T}$ for each task $k$, with the task-specific prompt parameterized as $\hat{P}_{k}=P^{*}\odot W_{k}$. The method combines multiple loss functions:
@@ 4.2 Reparameterization Methods
where the base model weights are quantized to 4-bit precision using NormalFloat (NF4), while the LoRA adapters ($L_{1},L_{2}$) operate in 16-bit precision. QLoRA integrates three major things that work together to enable efficient training: 4-bit NormalFloat (NF4), an information-theoretically optimal quantization format for normally distributed weights that outperforms standard 4-bit quantization; Double Quantization, which further compresses the model by quantizing the quantization constants themselves, reducing memory requirements by approximately 0.37 bits per parameter; and Paged Optimizers that manage memory spikes during training by utilizing NVIDIA unified memory for automatic CPU-GPU memory transfers. This combination enables fine-tuning of 65B parameter models on a single 48GB GPU while preserving full 16-bit performance. In practice, QLoRA achieves comparable results to full fine-tuning while requiring only 0.1% of the trainable parameters, demonstrating particular effectiveness in low-data regimes and better extrapolation to unseen topics.
@@
On the other hand, RoSA (Robust Adaptation) [103] combines both low-rank and sparse adapters trained in parallel with frozen pretrained weights. Inspired by robust principal component analysis, RoSA recognizes that while fine-tuning updates can be approximated by low-rank matrices [49], adding a sparse component can capture important outlier components that LoRA might miss. The method achieves comparable or better accuracy than full fine-tuning while maintaining the memory and computational efficiency of other PEFT methods, making it particularly effective for complex tasks like mathematical reasoning and coding where traditional LoRA methods often fall short. Additionally, RoSA is compatible with quantized base weights, making it the first approach to successfully combine quantization, low-rank projections, and sparsity while maintaining high performance.

### v2

@@ Abstract
The advent of large-scale pre-trained models has fundamentally transformed the field of artificial intelligence, unlocking new possibilities and achieving unprecedented performance across a wide range of tasks. Yet, these models inherit a fundamental limitation from traditional Machine Learning approaches: their heavy reliance on the i.i.d. assumption, which restricts their capacity to adapt to dynamic, real-world scenarios. We argue that the next major breakthrough in AI lies in developing systems capable of efficient and continual adaptation to evolving environments, where new data and tasks arrive sequentially. This need defines the domain of Continual Learning (CL), a Machine Learning paradigm aimed at developing neural models capable of learning throughout their lifespan without forgetting previous knowledge. In parallel, Parameter-Efficient Fine-Tuning (PEFT) methods have emerged as powerful tools for adapting large models to specific tasks with minimal computational resources. While PEFT techniques can match the performance of full fine-tuning with significantly fewer parameter updates, they remain vulnerable to catastrophic forgetting. This survey bridges the gap between CL and PEFT by focusing on the emerging field of Parameter-Efficient Continual Fine-Tuning (PECFT). We first provide a comprehensive overview of CL strategies and PEFT approaches, then we review recent advances in PECFT. We analyze existing methods, compare evaluation protocols and identify open challenges and promising directions for future research. Our aim is to underscore the synergy between CL and PEFT, offering insights and guidance to researchers seeking to build adaptive, scalable and efficient AI systems.
@@ 1 Introduction
Deep learning models achieve impressive results by scaling their parameters and the data used during training. However, these models often require substantial resources, which makes it impractical to update them continuously. Although Large Pretrained Models (PTMs) [24, 1, 94] demonstrate remarkable capabilities by leveraging extensive pretraining, which allows them to perform extremely well in specific domains. However, when deployed in real-world scenarios, fine-tuning these models becomes increasingly complex and computationally prohibitive, particularly as the number of tasks grows and data distributions continue to shift over time.
Consider a medical AI system that must continuously learn about new diseases, treatment protocols, and diagnostic techniques as medical knowledge advances. Traditional approaches would require either: (1) retraining the entire model from scratch, a computationally prohibitive and expensive process, or (2) fine-tuning on new data, which often leads to catastrophic forgetting of previously learned knowledge. Neither solution is practical for real-world deployment where systems must remain both current and comprehensive.
This challenge sits at the intersection of two fundamental limitations in current AI systems. First, while large pre-trained models excel at individual tasks, they struggle with sequential learning scenarios where new tasks arrive over time. Second, the enormous computational cost of updating these models makes frequent adaptation impractical, creating a tension between model capability and deployment efficiency.
To address these challenges, work in this area has focused on two complementary directions, Parameter Efficient Fine-Tuning (PEFT) and Continual Learning (CL). CL focuses on enabling models to learn from sequential tasks without forgetting previous knowledge, addressing the stability-plasticity dilemma that has long challenged machine learning systems, while PEFT offers techniques to adapt large pre-trained models using only a small subset of parameters, dramatically reducing computational costs while maintaining performance. The combination of these approaches gives rise to Parameter-Efficient Continual Fine-Tuning (PECFT), a new paradigm that combines the sequential learning capabilities of CL with the computational efficiency of PEFT. PECFT methods enable large pre-trained models to continuously evolve and adapt to new tasks while preserving past knowledge and maintaining computational efficiency. On one hand, CL enables models to continually adapt whiles PEFT ensures that efficiency whiles the model adapts.
Early works in this space, such as Learning to Prompt [153] and Dual Prompt [152], have demonstrated the potential of this convergence. These methods show that carefully designed parameter-efficient adaptations can enable continual learning without the traditional trade-offs between forgetting and computational cost. However, the field lacks a comprehensive understanding of the various approaches, their relative strengths and limitations, and the fundamental principles that govern effective PECFT.
The concept of PECFT represents more than just a technical advancement, it addresses a critical need for sustainable AI systems that can adapt and grow throughout their deployment lifecycle. Traditional machine learning paradigms assume static learning scenarios with fixed datasets and tasks. In contrast, real-world AI systems must operate in dynamic environments where:
  • 1.
    New tasks emerge continuously (e.g., new product categories in e-commerce, emerging medical conditions);
  • 2.
    Data distributions shift over time (e.g., changing user preferences, evolving social media trends);
  • 3.
    Computational resources are constrained (e.g., edge devices, cost-sensitive deployments);
  • 4.
    System downtime for retraining is unacceptable (e.g., autonomous vehicles, medical diagnosis systems).
PECFT offers a path toward AI systems that can meet these real-world demands by enabling efficient, incremental adaptation without compromising on performance or forgetting previous capabilities.
@@ 2.1.1 CL Challenges
Traditionally, machine learning models are usually trained on a finite, limited and static data set, limiting their adaptability. Naively training a model on new data to perform well on novel tasks can degrade its performance on previously learned tasks. This problem is known as catastrophic forgetting [108] and is the core challenge of CL. Theoretically, consequential re-training and adapting to new tasks can shift how features are represented within the model. This phenomenon, where the representation of features within a model evolves in a way that can negatively impact previous tasks, is known as representation drift [9, 57].
@@ 2.1.2 CL Scenarios
A second scenario is Class Incremental Learning (CIL) which, while similar to TIL in training, is more challenging in inference. This is because, during testing, the specific task information is unavailable, causing the model to have to predict among all observed classes, not only those of the task $t$. This lack of task information during testing requires the model to differentiate and classify all previously learned classes without additional context, making CIL significantly more challenging. Some variations of CIL include :
  1. 1.
@@ 3.1 Regularization-based CL
One alternative to naively training a model in a sequence of tasks is adding a constraint on the loss functions to encourage the mitigation of forgetting. This regularization can be imposed on weights by estimating the parameters importance so relevant weights do not drift significantly. One of the first approaches that proposed this was Elastic Weight Consolidation (EWC) [70], which captures the prior importance using a diagonally approximated Fisher information matrix.
Regularization can also be imposed on activations to prevent activation drift, which generally outperforms its weight-regularisation counterpart. One work in this line is Learning without Forgetting (LwF) [81], which prevents activations of the old network from drifting while learning new tasks. The less-forgetting learning penalizes the activation difference except for the fully-connected layer. Riemannian Walk [11] extends EWC by incorporating KL-divergence-based regularization, path integral methods to assess parameter importance and sample-based strategies for retaining past knowledge. Rotated EWC (R-EWC) [89] enhances EWC by reparameterizing the network to better align with the Fisher information matrix, improving the diagonal approximation and reducing forgetting.
@@ 3.4 Optimization-based CL
Table 2: Parameter-Efficient Fine-tuning Methods Overview
Method Key Mechanism and Features
Adapter-Based Methods
Adapter [49] Bottleneck layers (2kd/layer) with down/up projections. Effective task adaptation with minimal overhead
AdaptFormer [16] Specialized vision modules enhancing ViT transferability across tasks
Adapter Fusion [112] Two-stage fusion combining task adapters with knowledge distillation
Mera [44] Efficient integration of existing adapters using same-track combination strategy
AdaMix [150] Multiple parallel adapters providing diverse task perspectives
Invertible Adapters [113] Reversible bottleneck modules using NICE-style coupling for zero-shot cross-lingual transfer
Adapter+ [136] Lightweight variant with channel-wise scaling and optimized placement (only after multi-head attention)
Parallel Adapters [43] Run alongside feedforward sublayers rather than sequentially, enabling parallel computation
Mix-and-Match Adapters [43] Combine prefix tuning with scaled parallel adapters, achieving full fine-tuning performance
Prompt-Based Methods
Prompt Tuning [77] Continuous trainable prefix tokens (l×d) for task conditioning
P-Tuning [90] Hybrid discrete and continuous prompts with LSTM/MLP encoder
Prefix Tuning [79] Optimizable virtual tokens (0.1%) as prefix for generation
MPT [151] Distillation-based transferable prompts (0.035%) for multi-task scenarios
Reparameterization Methods
LoRA [52] Low-rank decomposition (2dr/layer) for weight updates. Zero inference overhead
QLoRA [22] 4-bit NormalFloat quantization + LoRA (0.1%) for memory-efficient training
DoRA [88] Separates magnitude and direction components for improved training dynamics
VB-LoRA [80] Constructs low-rank matrices from a shared vector bank
VeRA [71] Shared frozen matrices with learnable scaling vectors
Selective Methods
BitFit [165] Selective bias-term updates ($0.1\%$) preserving model knowledge
Masking [170] Task-specific binary masks identifying crucial weights (3-10%)
@@ 4 Parameter-Efficient Fine-Tuning
PEFT strategies can be broadly categorized based on how they modify or add parameters to the pre-trained model. These categories include: (i) additive methods, which introduce new task-specific parameters to the model; (ii) reparameterization methods, which reparameterize the existing model parameters more efficiently; (iii) and selective methods, which fine-tune a subset of the model parameters based on their importance to the task. Since this paper focuses on the intersection of Continual Learning and Parameter-Efficient Fine-tuning, we will present an overview of these methods, introducing and discussing only a few of these , for a comprehensive comparison of different PEFT categories and methods, we refer readers to recent surveys by [172] and [158]. A summary of the PEFT methods we discuss can be seen in Table 2. Following the overview of PEFT techniques, we analyze the computational efficiency of each approach, focusing on the number of trainable parameters introduces across various backbone models. To ensure a broad and comprehensive comparison, we include models from both NLP and Computer Vision domains, as well as a multi-modal backbone. This comparison, presented in Table 3, offers valuable insights for practitioners, providing efficiency metrics aside method descriptions to facilitate a clearer understanding of each PEFT strategy.
@@ 4.1.1 Adapter Based PEFT Methods
Invertible adapters [113] introduce a reversible bottleneck module inside the transformer. They use a NICE-style coupling mechanism: the input embedding is split in half and transformed by projections G, making the process invertible. This means no information is lost. They train these adapters with masked language modeling on unlabeled language data. In the MAD‑X framework, they stack invertible adapters along with language and task adapters. At inference, the English adapter is swapped out with a target-language one. This method keeps the base model frozen, adds only a small number of parameters, and supports zero-shot cross-lingual transfer—even to languages unseen during pre‑training.
Adapter+ [136] presents a lightweight yet robust variant that refines classic adapter designs through careful empirical analysis. Unlike standard approaches such as [50] Adapter+ places adapters only after multi-head attention, introduces channel-wise scaling, and adopts Houlsby-style [50] initialization. These adjustments yield strong out-of-the-box performance with minimal tuning, rivaling other adapter based methods.
Parallel adapters [43] differ from standard sequential adapters by running alongside the feedforward sublayer rather than after it. The authors add the adapter output directly to the base model’s output. This structure lets the adapter learn only residual corrections. It’s more flexible and can improve performance in transfer settings. It also enables parallel computation paths for faster inference.
Mix‑and‑Match adapters [43] combine beneficial features from various adapter approaches. They use prefix tuning at attention layers with a small bottleneck and scaled parallel adapters for feedforward layers. In their unified framework, adapters from different tasks can be combined. This variant matches full fine-tuning performance on some benchmarks while fine‑tuning only about 6.7% of parameters. Although it is a predominantly adapter-based, but they incorporate prompt-based ideas, specifically from [79].
@@ 4.1.2 Prompt Based PEFT Methods
Building on this foundation, P-Tuning [91] was introduced to address a critical limitation , the inherent instability issues of discrete prompts. It introduces trainable continuous prompt embeddings that are concatenated with traditional discrete prompts. The core template structure is defined as:
@@
where $P_{i}]$ represents continuous prompt embeddings that are optimized through backpropagation, while $x$ and $y$ represent the input and label respectively. The combination of trainable continuous embeddings with discrete prompts is done via a prompt encoder, i.e. LSTM or MLP, to model the dependencies between the embeddings. This hybrid approach significantly reduces performance variance, maintaining a consistent behavior where traditional methods showed up to 20% drops from single word changes in prompts.
@@
Taking prompt tuning to a multi-task setting, Multitask Prompt Tuning (MPT) [151] innovates by learning a single transferable prompt from multiple source tasks that can be adapted to target tasks. By operating in two stages , source training and target adaptation , MPT learns a task-shared prompt matrix through knowledge distillation. This prompt is decomposed into a task-shared matrix $P^{*}$ and low-rank task-specific matrices $W_{k}=u_{k}\otimes v_{k}^{T}$ for each task $k$, with the task-specific prompt parameterized as $\hat{P}_{k}=P^{*}\odot W_{k}$. The method combines multiple loss functions:
@@ 4.2 Reparameterization Methods
where the base model weights are quantized to 4-bit precision using NormalFloat (NF4), while the LoRA adapters ($L_{1},L_{2}$) operate in 16-bit precision. QLoRA integrates three major things that work together to enable efficient training: NF4, an information-theoretically optimal quantization format for normally distributed weights that outperforms standard 4-bit quantization; Double Quantization, which further compresses the model by quantizing the quantization constants themselves, reducing memory requirements by approximately 0.37 bits per parameter; and Paged Optimizers that manage memory spikes during training by utilizing NVIDIA unified memory for automatic CPU-GPU memory transfers. This combination enables fine-tuning of 65B parameter models on a single 48GB GPU while preserving full 16-bit performance. In practice, QLoRA achieves comparable results to full fine-tuning while requiring only 0.1% of the trainable parameters, demonstrating particular effectiveness in low-data regimes and better extrapolation to unseen topics.
@@
On the other hand, VB-LoRA [80] replaces traditional per-layer adapters with a global pool of reusable vectors, called a vector bank. Rather than learning unique low-rank matrices for each layer, VB-LoRA dynamically constructs them by selecting and combining vectors from this shared bank using a differentiable top-k selection mechanism. This design allows for extreme parameter sharing across the model, significantly reducing memory and storage overhead. Despite its compactness—requiring just a fraction of the parameters.
@@
Table 3: Parameter efficiency comparison of PEFT methods across different pre-trained model architectures. The table shows the number of trainable parameters and the percentage of total parameters updated for each method when applied to various foundation models
Model (Total Params) Method Trainable Params % Trainable
BERT Base Uncased [1.5mm](109B) Prompt [77] 15,360 0.01%
Prefix [79] 368,640 0.34%
LoRA [52] 443,906 0.41%
VeRA [71] 57,986 0.05%
DoRA [88] 471,554 0.43%
VB-LoRA [80] 509,442 0.47%
QLoRA [22] 443,906 0.40%
BitFit [165] 102,914 0.09%
ALBERT Base V2 [1.5mm](11B) Prompt [77] 15,360 0.13%
Prefix [79] 368,640 3.15%
LoRA [52] 38,402 0.33%
VeRA [71] 6,242 0.05%
DoRA [88] 40,706 0.35%
VB-LoRA [80] 103,938 0.89%
QLoRA [22] 38,402 0.32%
BitFit [165] 10,114 0.09%
RoBERTa Base [1.5mm](124B) Prompt [77] 15,360 0.01%
Prefix [79] 368,640 0.30%
LoRA [52] 1,034,498 0.83%
VeRA [71] 648,578 0.52%
DoRA [88] 1,062,146 0.85%
VB-LoRA [80] 1,100,034 0.88%
QLoRA [22] 1,034,498 0.82%
BitFit [165] 102,914 0.08%
ViT B16 [1.5mm](85B) Prompt [77] 15,360 0.01%
Prefix [79] 384,000 0.44%
LoRA [52] 442,368 0.52%
VeRA [71] 27,936 0.03%
DoRA[88] 470,016 0.55%
VB-LoRA[80] 802,816 0.54 %
QLoRA[22] 43,906 0.51%
Sequential adapter [49] 903,744 1.05 %
Double adapter [113] 451,872 0.05 %
Parallel adapter [42] 451,872 0.05 %
Adapter+ [136] 461,088 0.05 %
BitFit [165] 102,914 0.12%
CLIP ViT B16 [1.5mm](149B) Prompt [77] 15,360 0.01%
Prefix [79] 675,840 0.45%
LoRA [52] 737,280 0.49%
VeRA [71] 46,656 0.03%
DoRA [88] 783,360 0.52%
VB-LoRA [80] 802,816 0.54 %
QLoRA [88] 737,280 0.49%
Sequential adapter [49] 1,309,632 0.87 %
Double adapter [113] 654,816 0.43 %
Parallel adapter [42] 654,816 0.43 %
Adapter+ [136] 670,176 0.44 %
BitFit [165] 171,008 0.11%